# Deep Learning for 3D Human Pose Estimation and Mesh Recovery: A Survey

### v1

@@ 7 Applications
Motion Retargeting: Human motion retargeting can transfer human actions onto actors. Using pose estimation eliminates the need for motion capture systems and achieves image-to-image translation. Therefore, 3D human pose estimation is crucial for retargeting. Recently, there has been a significant amount of retargeting work based on 3D human pose estimation aberman2020skeleton ; yang2020transmomo ; yu2023bidirectionally . End-to-end methods and related datasets are also designed gomes2021shape . Additionally, unsupervised methods in In-the-Wild scenarios zhu2022mocanet have been developed, achieved through canonicalization operations and derived regularizations. Beyond bodily retargeting, facial retargeting mo2022towards ; chen2024morphable has also gained prominence, wherein the intricacies of facial expressions offer a refined portrayal of the actor’s emotional and psychological states.
@@
Action Recognition: Action recognition employs algorithms to identify and analyze human movements from images or videos. The results derived from 3D human pose estimation and reconstruction play a pivotal role in deciphering the dynamics of human motion within a three-dimensional context, thereby transforming these movements into actionable behavioral insights yang2021feedback ; mazzia2022action ; duan2022revisiting ; lu2023hard . Moreover, these advancements can significantly augment the efficiency of action recognition bian2021structural . It is also possible to address pose estimation and action recognition within the same framework through multi-task learning and sharing features luvizon2020multi .
@@
Security Monitoring: In video surveillance systems at public places or critical facilities, pedestrian tracking and re-identification are key tasks. Combined with human pose estimation, Human tracking is utilized for the surveillance and analysis of pedestrian flow, tracking specific targets, and tracking and analyzing human behavior in space. This approach is highly beneficial for tracking humans in complex scenarios bao2020pose ; reddy2021tessetrack ; goel2023humans ; sun2023trace . Considering the constrained viewpoints of individual cameras in such systems, Re-identification enables the recognition and tracking of the same person as they move across different camera views. Human pose estimation significantly contributes to the enhancement of this re-identification wang2021horeid .
@@
SLAM: SLAM precisely estimates its location by gathering sensor data from its environment, employing technologies such as cameras and LiDAR. It simultaneously constructs or refines an environmental map, facilitating self-localization and map creation in unfamiliar territories. Diverging from the conventional SLAM systems that predominantly concentrate on objects, recent advancements have shifted focus towards incorporating humans within the environmental context. Notably, Dai et al. dai2021indoor have illustrated the significance of 3D human trajectory reconstruction in indoor settings, enhancing indoor navigation capabilities. Furthermore, Kocabas et al. kocabas2023pace have innovatively integrated human motion priors into SLAM, effectively merging human pose estimation with scene analysis.
@@
Autonomous Driving: In robotic navigation and autonomous vehicle applications, estimating human pose enables these systems to comprehend human behavior and intentions better. This understanding facilitates more intelligent decision-making and interaction. Zheng et al. zheng2022multi propose a multi-modal approach employing 2D labels on RGB images as weak supervision for 3D human pose estimation in the context of autonomous vehicles. Wang et al. wang2023learning have developed a comprehensive framework to learn physically plausible human dynamics from real driving scenarios, effectively bridging the gap between actual and simulated human behavior in safety-critical applications.
@@
Human–Computer Interaction: Human-computer interaction entails the bidirectional communication between humans and computers, underscored by the critical need for computers to interpret human poses accurately. Liu et al. liu2022arhpe advanced this field by proposing asymmetric relation-aware representation learning for head pose estimation in industrial human–computer interaction, which utilizes an effective Lorentz distribution learning scheme. In Augmented Reality (AR) systems, the accuracy of human pose estimation significantly elevates the interaction quality between virtual objects and the real environment, fostering more natural interactions between humans and virtual entities. In this context, Weng et al. weng2019photo developed a novel method and application for animating a human subject from a single photo within AR.
@@ 8 Challenges and Conclusion
Large Models. The efficacy of large models in language anil2023palm ; achiam2023gpt and foundational computer vision tasks, such as segmentation kirillov2023segment and tracking yang2023track , has been universally acknowledged and is quite remarkable. Additionally, while there is burgeoning research in human-centric computer vision tasks ci2023unihcp , the area of 3D tasks still demands further investigation. Furthermore, the development of large models not only constitutes a significant area of research but also the exploration of their effective utilization presents substantial challenges and practical importance. Exploratory efforts to fuse pose estimation with large language models feng2023posegpt and the combination of large visual models with pose estimation are also meaningful.
More Detailed Reconstruction. At present, explicit model-based methodologies such as SMPL loper2015smpl and SMPL-X cai2023smpler , fall short of meeting the demands for detail that people expect. On the other hand, implicit representation approaches saito2019pifu ; huang2020arch ; he2020geo , as well as rendering techniques such as NeRF gao2022mps ; mu2023actorsnerf and Gaussian Splatting liu2023animatable , are capable of capturing fine details but lack sufficient robustness in pose estimation. Bridging the gap between robust pose estimation and surface details remains a formidable challenge that necessitates collaborative efforts from computer vision and computer graphics researchers.
Crowding and Occlusion Challenges. In open-world scenarios, the phenomena of crowding and occlusion are prevalent, and they represent long-standing challenges in the field of object detection. Currently, top-down methods depend on object detection, rendering these issues inescapable. Although bottom-up strategies may circumvent object detection, they confront formidable challenges regarding key point assembly.
@@
Speed. Speed is an essential aspect to consider in the practical deployment of algorithms. While most current research papers report achieving real-time performance on GPUs, a wide array of applications necessitates real-time and efficient processing on edge computing platforms, notably on ARM processors within smartphones. The disparity in performance between ARM processors and GPUs is pronounced, underscoring the immense value of optimizing for speed.

### v2

@@ 7 Applications
Motion Retargeting. Human motion retargeting can transfer human actions onto actors. Using pose estimation eliminates the need for motion capture systems and achieves image-to-image translation. Therefore, 3D human pose estimation is crucial for retargeting. Recently, there has been a significant amount of retargeting work based on 3D human pose estimation [253, 254, 255]. End-to-end methods and related datasets are also designed [256]. Additionally, unsupervised methods in In-the-Wild scenarios [257] have been developed, achieved through canonicalization operations and derived regularizations. Beyond bodily retargeting, facial retargeting [258, 259] has also gained prominence, wherein the intricacies of facial expressions offer a refined portrayal of the actor’s emotional and psychological states.
Human Avatars. Human avatars are human-like virtual entities created through digital technology, which can interact and express themselves on various digital media platforms. Human pose estimation and mesh recovery enable digital avatars to simulate human movements and behaviors more realistically and vividly, allowing developers to create virtual characters with behaviors highly similar to real individuals [209, 260]. For instance, Luo et al. [261] proposed a real-time multi-person avatar controller using noisy poses from video-based pose estimators and language-based motion generators.
@@
Action Recognition. Action recognition employs algorithms to identify and analyze human movements from images or videos. The results derived from 3D human pose estimation and reconstruction play a pivotal role in deciphering the dynamics of human motion within a three-dimensional context, thereby transforming these movements into actionable behavioral insights [262, 263, 1, 264]. Moreover, these advancements can significantly augment the efficiency of action recognition [265]. It is also possible to address pose estimation and action recognition within the same framework through multi-task learning and sharing features [266].
@@
Security Monitoring. In video surveillance systems at public places or critical facilities, pedestrian tracking and re-identification are key tasks. Combined with human pose estimation, Human tracking is utilized for the surveillance and analysis of pedestrian flow, tracking specific targets, and tracking and analyzing human behavior in space. This approach is highly beneficial for tracking humans in complex scenarios [267, 268, 269, 270]. Considering the constrained viewpoints of individual cameras in such systems, Re-identification enables the recognition and tracking of the same person as they move across different camera views. Human pose estimation significantly contributes to the enhancement of this re-identification [11].
@@
SLAM. SLAM precisely estimates its location by gathering sensor data from its environment, employing technologies such as cameras and LiDAR. It simultaneously constructs or refines an environmental map, facilitating self-localization and map creation in unfamiliar territories. Diverging from the conventional SLAM systems that predominantly concentrate on objects, recent advancements have shifted focus towards incorporating humans within the environmental context. Notably, Dai et al. [271] have illustrated the significance of 3D human trajectory reconstruction in indoor settings, enhancing indoor navigation capabilities. Furthermore, Kocabas et al. [272] have innovatively integrated human motion priors into SLAM, effectively merging human pose estimation with scene analysis.
@@
Autonomous Driving. In robotic navigation and autonomous vehicle applications, estimating human pose enables these systems to comprehend human behavior and intentions better. This understanding facilitates more intelligent decision-making and interaction. Zheng et al. [14] propose a multi-modal approach employing 2D labels on RGB images as weak supervision for 3D human pose estimation in the context of autonomous vehicles. Wang et al. [15] have developed a comprehensive framework to learn physically plausible human dynamics from real driving scenarios, effectively bridging the gap between actual and simulated human behavior in safety-critical applications.
@@
Human–Computer Interaction. Human-computer interaction entails the bidirectional communication between humans and computers, underscored by the critical need for computers to interpret human poses accurately. Liu et al. [12] advanced this field by proposing asymmetric relation-aware representation learning for head pose estimation in industrial human–computer interaction, which utilizes an effective Lorentz distribution learning scheme. In Augmented Reality (AR) systems, the accuracy of human pose estimation significantly elevates the interaction quality between virtual objects and the real environment, fostering more natural interactions between humans and virtual entities. In this context, Weng et al. [16] developed a novel method and application for animating a human subject from a single photo within AR.
@@ 8 Challenges and Conclusion
Speed. Speed is an essential aspect to consider in the practical deployment of algorithms. While most current research papers report achieving real-time performance on GPUs, a wide array of applications necessitates real-time and efficient processing on edge computing platforms, notably on ARM processors within smartphones. The disparity in performance between ARM processors and GPUs is pronounced, underscoring the immense value of optimizing for speed. Although some video-based visual processing tasks [273, 274, 275] have already acknowledged the importance of processing speed for real-time application scenarios, further research is still needed in the context of 3D cases [77].
Crowding and Occlusion Challenges. In open-world scenarios, the phenomena of crowding and occlusion are prevalent, and they represent long-standing challenges in the field of object detection. Currently, top-down methods depend on object detection, rendering these issues inescapable. Although bottom-up strategies may circumvent object detection, they confront formidable challenges regarding key point assembly. Therefore, some researchers [121] attempt to integrate both methods together. Moreover, incorporating temporal consistency constraints (such as optical flow [102]) during mining can also alleviate occlusion issues.
Large Models. The efficacy of large models in language [276, 277] and foundational computer vision tasks, such as segmentation [278] and tracking [279], has been universally acknowledged and is quite remarkable. Additionally, while there is burgeoning research in human-centric computer vision tasks [280], the area of 3D tasks still demands further investigation. Furthermore, the development of large models not only constitutes a significant area of research but also the exploration of their effective utilization presents substantial challenges and practical importance. Exploratory efforts to fuse pose estimation with large language models [281] and the combination of large visual models with pose estimation are also meaningful.
More Detailed Reconstruction. At present, explicit model-based methodologies such as SMPL [35] and SMPL-X [61], fall short of meeting the demands for detail that people expect. On the other hand, implicit representation approaches [193, 195, 198], as well as rendering techniques such as NeRF [201, 51] and Gaussian Splatting [211], are capable of capturing fine details but lack sufficient robustness in pose estimation. Bridging the gap between robust pose estimation and surface details remains a formidable challenge that necessitates collaborative efforts from computer vision and computer graphics researchers.
Reconstructing with the Environment. Humans engage with various objects and environments within the real world in the daily life. The ability to reconstruct these interactions in 3D is paramount for comprehending and simulating their dynamics. This aspect holds particular significance in the realm of artificial intelligence applied to human avatars. In the reconstruction of object interactions, challenges such as pose initialization and sparse observation of objects in Structure from Motion are encountered [7]. Moreover, Yi et al. [282] indicate a reciprocal enhancement between environmental and human reconstruction, resulting in enhanced effectiveness across both domains.
Controllability and Animatability. Controllability refers to the precise control over the postures, movements, and expressions of human avatars, enabling them to exhibit various desired behaviors in virtual environments. Animation, on the other hand, denotes the ability of avatars to vividly and smoothly demonstrate various movements and expressions, making them appear more realistic and natural. Jiang et al. [283] employed an explicit deformation field to deform the neural implicit field, thus enabling the animation of human characters. They mapped the target human body mesh to the template human body mesh, represented by a parameterized human body model, thereby simplifying the animation and reshaping of the generated avatars through the control of pose and shape parameters. By enhancing the controllability and animatability of human body mesh reconstruction, developers can create digital characters that are more realistic, vivid, and expressive. This, in turn, enhances their interactivity and appeal in the virtual world.